OpenStudy (anonymous):

Probability: A fair six-sided die is tossed. What is the probability that a number less than 3 will show on the face of the die? a. 1/2 b. 1/3 c. 2/3 d. 1/6

OpenStudy (ajprincess):

What is the probability of getting 3?

OpenStudy (ajprincess):

@rosimary

OpenStudy (anonymous):

the probability of getting anything less than a three would be half, minus one..so B;)

OpenStudy (anonymous):

i dont think that made sense..

OpenStudy (anonymous):

oops! not B i ment A;) sry!

OpenStudy (ajprincess):

There are two numbers that are less than 3 in a dice. Total number of numbers available is 6. So ur probability will be 2/6=1/3 Am I clear @rosimary?

OpenStudy (anonymous):

because there is only 2 sides that the dice could land on instead of 6

OpenStudy (anonymous):

wait.. can the dice land on the three? if it can, then yes, the answer is 1/3 but if it can only be less than a three then the answer is 1/2

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

so the answer would be 1/2 because it has to be any number less than 3 which is 2

OpenStudy (anonymous):

and 1.. so only 2 #'s

OpenStudy (anonymous):

@rosimary does this explanation of dice-rolling help?

OpenStudy (anonymous):

ok wait I'm confused. I was at first understanding the dice thing but as I kept reading the different answers I got confused

OpenStudy (anonymous):

Let's start over... @ajprincess has a great explanation.

OpenStudy (anonymous):

sry 4 confusing:/

OpenStudy (anonymous):

@kelly226 was right too... :) Just lots more to read. Let's have one person explain :) and make it easier for @rosimary to read.

OpenStudy (anonymous):

good idea;)

OpenStudy (anonymous):

Ok, I'll go... are you with me, @rosimary ?

OpenStudy (anonymous):

When you roll a die (dice?), you can get any one of six outcomes... the numbers 1-6, right? So, the chance of getting any specific number, for example, rolling a 4, is the number of ways you can get that 4 (just one way... roll a 4!!) DIVIDED by the number of possible rolls (which is 6, right?) So the chance of getting a 4 is 1 / 6 = (1/6). Your problem asks you for the chance of getting a number less than 3. There are two numbers less than 3.... you could roll a 1 or you could roll a 2. So two different rolls work. There are still 6 possible numbers you COULD roll... 1,2,3,4,5, or 6. But only rolls of 1 or 2 work. The chance of rolling a number less than 3 is the number of ways that work (2) divided by the total number of possible outcomes (6). So it's 2/6 which simplifies to 1/3

OpenStudy (anonymous):

@rosimary please take your time and read that explanation and see if it helps. I think you probably almost had it earlier and then got confused the more you thought about it.

OpenStudy (anonymous):

Got it. Thanks for the explanation. I also have a writing question opened if anyone could help I would greatly appreciate it.

OpenStudy (anonymous):

that's not my area of expertise... sorry!! Probably should post it as a new question, preferably in the writing area. Good luck!